OpenStudy (anonymous):

Margin of error Surveys conducted in American high schools concluded that 90% of the students in a sample of 400 students had more than one active email account. What would the margin of error be for the population proportion? 1.5% 3% 4.5% 9% I think 1.5% but just checking. @perl

OpenStudy (perl):

are you given a signficance level?

OpenStudy (anonymous):

no

OpenStudy (perl):

we can plug in signficance levels. lets start with 95% significance

OpenStudy (perl):

The margin of error for a sample proportion is Zc*SE . Zc is the critical z score , which depends on level of significance. SE = sqrt(p*q/n) where p is the population proportion q is 1-p, and n is the sample size. Typically, for a confodence interval, the sample proportion and its complement are substituted for p and q.)

OpenStudy (perl):

Margin of Error = Zc * √(.90 * .10 / 400 )

OpenStudy (anonymous):

I got 0.015. What is the Zc?

OpenStudy (perl):

Margin of Error = Zc * √(.90 * .10 / 400 ) Zc is the critical z value the higher significance you want, the higher the z score (the bigger the margin of error)

OpenStudy (perl):